# Non-resonant relaxation of rotating globular clusters

### v2

@@
The long-term relaxation of rotating, spherically symmetric globular clusters is investigated through an extension of the orbit averaged Chandrasekhar non-resonant formalism. A comparison is made with the long-term evolution of the distribution function in action space, measured from averages of sets of $N$-body simulations up to core collapse. The impact of rotation on in-plane relaxation is found to be weak. In addition, we observe a clear match between theoretical predictions and $N$-body measurements. For the class of rotating models considered, we find no strong gravo-gyro catastrophe accelerating core collapse. Both kinetic theory and simulations predict a reshuffling of orbital inclinations from overpopulated regions to underpopulated ones. This trend accelerates as the amount of rotation is increased. Yet, for orbits closer to the rotational plane, the non-resonant prediction does not reproduce numerical measurements. We argue that this mismatch stems from these orbits’ coherent interactions, which are not captured by the non-resonant formalism that only addresses local deflections.
@@ 1 Introduction
Rotation is ubiquitous in stellar systems. In effect, it provides a source of free energy, allowing clusters to efficiently reshuffle their orbital structure towards more likely configurations. Yet, historically, the study of globular clusters has been mainly focused on isotropic, non-rotating, old globular clusters (Aarseth et al., 1974; Spitzer, 1975; Cohn, 1979; Trager et al., 1995; Miocchi et al., 2013). The reasons behind such simplifications are two-fold: (i) naturally, it is easier numerically and analytically to neglect the effect of rotation; (ii) spherical isotropic models – e.g., the King models (King, 1966) or the Wilson models (Wilson, 1975) – provided a satisfactory zeroth-order description of the main observed dynamical properties of globular clusters (see, e.g., McLaughlin & van der Marel, 2005).
@@
The last decade has seen the extraction of new data – e.g., HST (Bellini et al., 2017) and Gaia DR2 (Bianchini et al., 2018; Sollima et al., 2019). These surveys gave the astrophysical community access to numerous and detailed observations of the internal kinematics of several globular clusters of the Milky Way (Bianchini et al., 2013; Fabricius et al., 2014; Watkins et al., 2015; Ferraro et al., 2018; Kamann et al., 2018), as well as a quantification of the degree of velocity anisotropy (Jindal et al., 2019). Using these new data sets, the historical highly symmetric cluster models are not satisfactory anymore.
@@
Therefore, a secular theory which describes the evolution of globular clusters and accounts for their rotation is needed to describe their long-term evolution. In particular, the Fokker-Planck theory was used to probe the impact of rotation of the long-term evolution of such clusters, both before (Einsel & Spurzem, 1999) and after (Kim et al., 2002) core collapse. Furthermore, observations show that the angular momentum distribution measured in Galactic clusters retain the signature of their formation process (Lanzoni et al., 2018). While $N$-body simulations are able to reproduce these results (Tiongco et al., 2016, 2022), the historical context and the complexity of the problem have led to few analytical explorations (Geyer et al., 1983; White & Shawl, 1987; Kontizas et al., 1989). These remain scarce even now (see, e.g., Stetson et al., 2019; Rozier et al., 2019; Livernois et al., 2022).
@@
In isolated systems, a non-zero total angular momentum, i.e. the presence of rotation, can have a significant impact on the cluster’s long-term evolution, e.g., through the gravo-gyro catastrophe (Hachisu, 1979; Ernst et al., 2007). This phenomenon has been observed in a range of rotating systems, including gas cylinders (Inagaki & Hachisu, 1978; Hachisu, 1979), gaseous discs (Hachisu, 1982), flattened (quasi-spherical) star clusters (Akiyama & Sugimoto, 1989; Einsel & Spurzem, 1999; Ernst et al., 2007), and clusters with embedded black holes (Fiestas & Spurzem, 2010; Kamlah et al., 2022). Let us however note that many of these studies used a rotating King model to study the impact of rotation (see, e.g., Einsel & Spurzem, 1999; Varri & Bertin, 2012). Therein, changing the amount of rotation impacts the density profile: this makes comparisons amongst different models less clear.
@@
The concurrent occurrence of internal rotation and a spectrum of stellar masses can result in the formation of an oblate core of fast rotating heavy masses (Kim et al., 2004; Tiongco et al., 2021). More precisely, the orbital inclinations of the heaviest stars align w.r.t. one another, inducing a mass segregation in the distribution of orbital inclinations (Szölgyén et al., 2019). The generally agreed explanation for this phenomenon is resonant relaxation and resonant friction (Rauch & Tremaine, 1996; Meiron & Kocsis, 2019). While this effect concerns globular clusters, nuclear cluster with a dominant massive BH (Szölgyén & Kocsis, 2018; Foote et al., 2020; Gruzinov et al., 2020; Magnan et al., 2022; Ginat et al., 2023) also display this spontaneous alignment.
@@ 2 Non-resonant relaxation
The long-term evolution of the cluster’s DF is described by the Fokker–Planck (FP) equation (see, e.g., §7.4 of Binney & Tremaine, 2008)
@@ 2.2 Orbit average and secular evolution
where $X$ runs over all the diffusion coefficients to compute, i.e. runs over singlets and pairs from ${\{E,L,L_{z}\}}$. Here, $\Omega_{r}$ is the radial frequency and $v_{\mathrm{r}}$ the radial velocity that is evaluated along the mean field orbit. We refer to appendix C for detailed expressions. In particular, in section C.1, we show how the angular average can be performed explicitly for some classes of rotating models. And, in section C.2, we use the same approach as in appendix F2 of T22 to compute the radial average using a stable numerical integration scheme.
The final stage of the calculation is to convert the orbit averaged diffusion coefficients in $(E,L,L_{z})$ into ones in ${\bm{J}\!=\!(J_{\mathrm{r}},L,L_{z})}$. This is rather straightforward, as detailed in appendix D. Lastly, these diffusion coefficients are those used to evaluate the FP equation (1). Along the same lines, one can similarly describe the dynamics within the coordinates ${\bm{J}_{\mathrm{c}}\!=\!(J_{\mathrm{r}},L,\cos I)}$.
@@ 3 Long-term relaxation
First, let us describe the classes of clusters considered. The in-plane distribution of orbits, i.e. when projecting in the ${(J_{\mathrm{r}},L)}$ space, follows the same Plummer DFs (see, e.g., Dejonghe, 1987) as in section 3.1 of T22. Velocity anisotropy is encoded with the dimensionless parameter $q$, with $q=0$ corresponding to an isotropic distribution, and ${q\!>\!0}$ (resp. ${q\!<\!0}$) associated with radially (resp. tangentially) anisotropic velocity distributions.
@@ 3.1 Sphericity of the cluster
The NR theory presented in section 2 assumes that the cluster remains spherically symmetric throughout. However, Rozier et al. (2019) showed that (sufficiently) rotating clusters can harbour unstable modes (see, e.g., fig. 8 therein). To avoid this complication, we restrict ourselves to (linearly) stable rotating clusters, i.e. clusters that would remain spherically symmetric.
To probe the conservation of spherical symmetry, we investigate the sphericity, $h$, of clusters, as defined in appendix F in terms of the eigenvalue ratio of the density-weighted inertia tensor. This is illustrated in Fig. 2.
@@ 3.2 The gravo-gyro catastrophe
with $r_{\mathrm{h}}$ the half-mass radius. For a cluster with ${N\!=\!10^{5}}$ stars, this yields ${t_{\mathrm{rh}}\!=\!994\,{\mathrm{HU}}}$, where ${\mathrm{HU}}$ stands for the Hénon units (Heggie & Mathieu, 1986) in which ${G\!=\!M\!=\!R_{\mathrm{v}}\!=\!1}$, with $R_{\mathrm{v}}$ the virial radius [^3] For a Plummer cluster, this is related to the Plummer scale length $b$ by the relation ${R_{\mathrm{v}}\!=\!16b/(3\pi)}$ (see, e.g., Dejonghe, 1987).. In practice, with ${N\!=\!10^{5}}$ running a numerical simulation up to ${t\!\sim\!t_{\mathrm{rh}}}$ takes about one day of computation, see appendix E. In Breen et al. (2017), core collapse occurs at ${t\!\sim\!\!17t_{\mathrm{rh}}}$ for an isotropic cluster (see table 1 therein for the dependence w.r.t. velocity anisotropy). Hence, integrating clusters with ${N\!=\!10^{5}}$ up to core collapse is not reasonably feasible. In this section we therefore scale down the size of the clusters to ${N\!=\!10^{4}}$ stars. In that case, ${t_{\mathrm{rh}}\!=\!132\,{\mathrm{HU}}}$, and core collapse is numerically reached in about ten hours. Since we will focus on measuring the core radius – a very integrated quantity – the quality of the numerical measurements will not be too much degraded by this use of a smaller value for $N$. In practice, measurements are averaged over 50 runs.
@@ 3.3 In-plane vs out-of-plane diffusion
Using once again $N$-body simulations with ${N\!=\!10^{4}}$ stars, let us finally investigate the typical features of the in-plane diffusion, i.e. relaxation in ${(J_{\mathrm{r}},L)}$, and out-of-plane diffusion, i.e. relaxation in ${\cos I}$
First, Fig. 4 reports on the time evolution of the DF in $L$, i.e. the in-plane relaxation of the distribution of angular momenta.
@@
As could have been expected, the clusters tend towards smoother distribution of inclinations, i.e. relaxation tends to erase discontinuities. The sharp discontinuity at ${\cos I\!=\!0}$ has already been washed away after $1\,t_{\mathrm{rh}}$, regardless of the initial amplitude of the discontinuity. We refer to appendix K.2 for a comparison of the early relaxation between a discontinuous and a smooth distribution of orbital inclinations. Note that the redistribution of prograde and retrograde orbits observed in Fig. 5 does not conflict with the conservation of the cluster’s total angular momentum. Indeed, the average of ${\cos I\!=\!L_{z}/L}$ is not constrained by any global invariance, contrary to $L_{z}$. As such, even if the number of particles with positive and negative $\cos I$ (hence $L_{z}$) changes, the conservation of the total angular momentum is ensured by a modulation of the norm of each star’s angular momentum vector.
@@ 4 In-plane diffusion
Assuming that the cluster’s relaxation is driven by local pairwise deflections, its long-term evolution is described by the FP equation (1). While this equation formally describes the evolution in 3D action space, it is convenient to study relaxation in two-dimensional projections, in particular to improve the signal-to-noise ratio. Yet, such a projection comes at the cost of more intricate theoretical predictions, that require additional integrations along some third action. In this section, following the same approach as in T22, we first focus on “in-plane relaxation” i.e. relaxation occurring in ${(J_{\mathrm{r}},L)}$. Our goal is to compare quantitatively the $N$-body measurements with the NR prediction.
@@ 4.1 $N$-body measurements
As in Fig. 3, we recover that rotation does not really impact the rate of core collapse in these linearly stable clusters. Conversely, when the clusters are unstable, e.g., for ${(q,\alpha)\!=\!(1,0.5)}$ (see Fig. 2), we find numerically (not reported in Fig. 6) that the clusters flatten and that relaxation is accelerated.
@@ 4.2 Non-resonant prediction
Let us now compute the NR prediction for the relaxation rate in ${(J_{\mathrm{r}},L)}$. This prediction is numerically very intensive and requires the evaluation of five embedded integrals, as demonstrated in Appendices A and H. Importantly, we stress that in the present case, the LBD (equation 5) allows us to analytically perform the angular part of the orbit average (equation 4). This helps reducing the computation time. We refer to Appendix C.1 for details.
@@
Figure 8: Relaxation rate, $\partial F/\partial t$, in ${(J_{\mathrm{r}},L)}$ as in Fig. 7, but here from the NR theory. The prediction matches the $N$-body measurements from Fig. 7, up to an overall prefactor which depends on $q$, and (very) weakly on $\alpha$.
@@
When comparing the $N$-body measurements (Fig. 7) with the NR prediction (Fig. 8), the NR theory successfully recovers the in-plane relaxation of stable rotating clusters: both figures exhibit strikingly similar structures in action space. We obtain results for the non-rotating case which are in agreement with T22. Yet, even though amplitudes in both approaches are comparable, there is still a (slight) overall prefactor mismatch. The inclusion of rotation has little impact on this mismatch. A more quantitative comparison would require performing (many) more $N$-body runs, as well as a more precise measurement of the relaxation rate. This would be no light undertaking, and will be the subject of future works.
@@ 5.1 $N$-body measurements
Let us now consider the rotating clusters presented in the middle and bottom rows of Fig. 9. In radially anisotropic (${q\!=\!1}$) and isotropic (${q\!=\!0}$) clusters, the systems lose stars in the prograde region (${\cos I\!>\!0}$) and gain stars in the retrograde region (${\cos I\!<\!0}$). As one increases $\alpha$, i.e. as one increases the net rotation, this trend strengthens. This is in agreement with the orbital reshuffling observed in Fig. 8, where the clusters always tend to isotropise their in-plane distribution. Finally, we note that the highest diffusion rates in inclination are observed near ${\cos I\!=\!0}$. This corresponds precisely to the location of the discontinuity of the LBD.

### v3

@@
The long-term relaxation of rotating, spherically symmetric globular clusters is investigated through an extension of the orbit-averaged Chandrasekhar non-resonant formalism. A comparison is made with the long-term evolution of the distribution function in action space, measured from averages of sets of $N$-body simulations up to core collapse. The impact of rotation on in-plane relaxation is found to be weak. In addition, we observe a clear match between theoretical predictions and $N$-body measurements. For the class of rotating models considered, we find no strong gravo-gyro catastrophe accelerating core collapse. Both kinetic theory and simulations predict a reshuffling of orbital inclinations from overpopulated regions to underpopulated ones. This trend accelerates as the amount of rotation is increased. Yet, for orbits closer to the rotational plane, the non-resonant prediction does not reproduce numerical measurements. We argue that this mismatch stems from these orbits’ coherent interactions, which are not captured by the non-resonant formalism that only addresses local deflections.
@@ 1 Introduction
Rotation is ubiquitous in stellar systems. In effect, it provides a source of free energy, allowing clusters to efficiently reshuffle their orbital structure towards more likely configurations. Yet, historically, the study of globular clusters has been mainly focused on isotropic, non-rotating, old globular clusters (Aarseth et al., 1974; Spitzer, 1975; Cohn, 1979; Trager et al., 1995; Miocchi et al., 2013). The reasons behind such simplifications are two-fold: (i) naturally, it is easier numerically and analytically to neglect the effect of rotation; (ii) spherical isotropic models – for example, the King models (King, 1966) or the Wilson models (Wilson, 1975) – provided a satisfactory zeroth-order description of the main observed dynamical properties of globular clusters (see, e.g., McLaughlin & van der Marel, 2005).
@@
The last decade has seen the extraction of new data – for instance, HST (Bellini et al., 2017) and Gaia DR2 (Bianchini et al., 2018; Sollima et al., 2019). These surveys gave the astrophysical community access to numerous and detailed observations of the internal kinematics of several globular clusters of the Milky Way (Bianchini et al., 2013; Fabricius et al., 2014; Watkins et al., 2015; Ferraro et al., 2018; Kamann et al., 2018), as well as a quantification of the degree of velocity anisotropy (Jindal et al., 2019). Using these new data sets, the historical highly symmetric cluster models are not satisfactory anymore.
@@
Therefore, a secular theory that describes the evolution of globular clusters and accounts for their rotation is needed to describe their long-term evolution. In particular, the Fokker-Planck theory was used to probe the impact of rotation of the long-term evolution of such clusters, both before (Einsel & Spurzem, 1999) and after (Kim et al., 2002) core collapse. Furthermore, observations show that the angular momentum distribution measured in Galactic clusters retain the signature of their formation process (Lanzoni et al., 2018). While $N$-body simulations are able to reproduce these results (Tiongco et al., 2016, 2022), the historical context and the complexity of the problem have led to few analytical explorations (Geyer et al., 1983; White & Shawl, 1987; Kontizas et al., 1989). These remain scarce even now (see, e.g., Stetson et al., 2019; Rozier et al., 2019; Livernois et al., 2022).
@@
In isolated systems, a non-zero total angular momentum, that is, the presence of rotation, can have a significant impact on the cluster’s long-term evolution, for instance, through the “gravo-gyro catastrophe” (Hachisu, 1979; Ernst et al., 2007). This phenomenon has been observed in a range of rotating systems, including gas cylinders (Inagaki & Hachisu, 1978; Hachisu, 1979), gaseous discs (Hachisu, 1982), flattened (quasi-spherical) star clusters (Akiyama & Sugimoto, 1989; Einsel & Spurzem, 1999; Ernst et al., 2007), and clusters with embedded black holes (Fiestas & Spurzem, 2010; Kamlah et al., 2022). Let us however note that many of these studies used a rotating King model to study the impact of rotation (see, e.g., Einsel & Spurzem, 1999; Varri & Bertin, 2012). Therein, changing the amount of rotation impacts the density profile: this makes comparisons amongst different models less clear.
@@
The concurrent occurrence of internal rotation and a spectrum of stellar masses can result in the formation of an oblate core of fast rotating heavy masses (Kim et al., 2004; Tiongco et al., 2021). More precisely, the orbital inclinations of the heaviest stars align with respect to one another, inducing a mass segregation in the distribution of orbital inclinations (Szölgyén et al., 2019). The generally agreed explanation for this phenomenon is resonant relaxation and resonant friction (Rauch & Tremaine, 1996; Meiron & Kocsis, 2019). While this effect concerns globular clusters, nuclear cluster with a dominant massive BH (Szölgyén & Kocsis, 2018; Foote et al., 2020; Gruzinov et al., 2020; Magnan et al., 2022; Ginat et al., 2023) also display this spontaneous alignment.
@@ 2 Non-resonant relaxation
The long-term evolution of the cluster’s DF is described by the Fokker–Planck (FP) equation (see, e.g., section 7.4 of Binney & Tremaine, 2008)
@@ 2.2 Orbit average and secular evolution
where $X$ runs over all the diffusion coefficients to compute, that is, runs over singlets and pairs from ${\{E,L,L_{z}\}}$. Here, $\Omega_{r}$ is the radial frequency and $v_{\mathrm{r}}$ the radial velocity that is evaluated along the mean field orbit. We refer to appendix C for detailed expressions. In particular, in appendix C.1, we show how the angular average can be performed explicitly for some classes of rotating models. And, in appendix C.2, we use the same approach as in appendix F2 of T22 to compute the radial average using a stable numerical integration scheme.
The final stage of the calculation is to convert the orbit-averaged diffusion coefficients in $(E,L,L_{z})$ into ones in ${\bm{J}\!=\!(J_{\mathrm{r}},L,L_{z})}$. This is rather straightforward, as detailed in appendix D. Lastly, these diffusion coefficients are those used to evaluate the FP equation (1). Along the same lines, one can similarly describe the dynamics within the coordinates ${\bm{J}_{\mathrm{c}}\!=\!(J_{\mathrm{r}},L,\cos I)}$.
@@ 3 Long-term relaxation
First, let us describe the classes of clusters considered. The in-plane distribution of orbits (i.e., when projecting in the ${(J_{\mathrm{r}},L)}$ space) follows the same Plummer DFs (see, e.g., Dejonghe, 1987) as in section 3.1 of T22. Velocity anisotropy is encoded with the dimensionless parameter $q$, with $q=0$ corresponding to an isotropic distribution, and ${q\!>\!0}$ (resp. ${q\!<\!0}$) associated with radially (resp. tangentially) anisotropic velocity distributions.
@@ 3.1 Sphericity of the cluster
The NR theory presented in section 2 assumes that the cluster remains spherically symmetric throughout. However, Rozier et al. (2019) showed that (sufficiently) rotating clusters can harbour unstable modes (see, e.g., fig. 8 therein). To avoid this complication, we restricted ourselves to (linearly) stable rotating clusters, that is, clusters that would remain spherically symmetric.
To probe the conservation of spherical symmetry, we investigateed the “sphericity”, $h$, of clusters, as defined in appendix F in terms of the eigenvalue ratio of the density-weighted inertia tensor. This is illustrated in Fig. 2.
@@ 3.2 The gravo-gyro catastrophe
with $r_{\mathrm{h}}$ the half-mass radius. For a cluster with ${N\!=\!10^{5}}$ stars, this yields ${t_{\mathrm{rh}}\!=\!994\,{\mathrm{HU}}}$, where ${\mathrm{HU}}$ stands for the Hénon units (Heggie & Mathieu, 1986) in which ${G\!=\!M\!=\!R_{\mathrm{v}}\!=\!1}$, with $R_{\mathrm{v}}$ the virial radius [^3] For a Plummer cluster, this is related to the Plummer scale length $b$ by the relation ${R_{\mathrm{v}}\!=\!16b/(3\pi)}$ (see, e.g., Dejonghe, 1987).. In practice, with ${N\!=\!10^{5}}$ running a numerical simulation up to ${t\!\sim\!t_{\mathrm{rh}}}$ took about one day of computation, see appendix E. In Breen et al. (2017), core collapse occurs at ${t\!\sim\!\!17t_{\mathrm{rh}}}$ for an isotropic cluster (see table 1 therein for the dependence with respect to velocity anisotropy). Hence, integrating clusters with ${N\!=\!10^{5}}$ up to core collapse is not reasonably feasible. In this section we therefore scaled down the size of the clusters to ${N\!=\!10^{4}}$ stars. In that case, ${t_{\mathrm{rh}}\!=\!132\,{\mathrm{HU}}}$, and core collapse was numerically reached in about ten hours. Since we focused on measuring the core radius – a very integrated quantity – the quality of the numerical measurements was not too much degraded by this use of a smaller value for $N$. In practice, measurements were averaged over 50 runs.
@@ 3.3 In-plane vs out-of-plane diffusion
Using once again $N$-body simulations with ${N\!=\!10^{4}}$ stars, let us finally investigate the typical features of the in-plane diffusion (i.e. relaxation in ${(J_{\mathrm{r}},L)}$) and out-of-plane diffusion (i.e. relaxation in ${\cos I}$).
First, Fig. 4 reports on the time evolution of the DF in $L$, that is, the in-plane relaxation of the distribution of angular momenta.
@@
As could have been expected, the clusters tend towards smoother distribution of inclinations, meaning that relaxation tends to erase discontinuities. The sharp discontinuity at ${\cos I\!=\!0}$ has already been washed away after $1\,t_{\mathrm{rh}}$, regardless of the initial amplitude of the discontinuity. We refer to appendix K.2 for a comparison of the early relaxation between a discontinuous and a smooth distribution of orbital inclinations. Note that the redistribution of prograde and retrograde orbits observed in Fig. 5 does not conflict with the conservation of the cluster’s total angular momentum. Indeed, the average of ${\cos I\!=\!L_{z}/L}$ is not constrained by any global invariance, contrary to $L_{z}$. As such, even if the number of particles with positive and negative $\cos I$ (hence $L_{z}$) changes, the conservation of the total angular momentum is ensured by a modulation of the norm of each star’s angular momentum vector.
@@ 4 In-plane diffusion
Assuming that the cluster’s relaxation is driven by local pairwise deflections, its long-term evolution is described by the FP equation (1). While this equation formally describes the evolution in 3D action space, it is convenient to study relaxation in two-dimensional projections, in particular to improve the signal-to-noise ratio. Yet, such a projection comes at the cost of more intricate theoretical predictions, that require additional integrations along some third action. In this section, following the same approach as in T22, we first focus on in-plane relaxation, that is, relaxation occurring in ${(J_{\mathrm{r}},L)}$. Our goal is to compare quantitatively the $N$-body measurements with the NR prediction.
@@ 4.1 $N$-body measurements
As in Fig. 3, we recover that rotation does not really impact the rate of core collapse in these linearly stable clusters. Conversely, when the clusters are unstable, such as for ${(q,\alpha)\!=\!(1,0.5)}$ (see Fig. 2), we find numerically (not reported in Fig. 6) that the clusters flatten and that relaxation is accelerated.
@@ 4.2 Non-resonant prediction
Let us now compute the NR prediction for the relaxation rate in ${(J_{\mathrm{r}},L)}$. This prediction was numerically very intensive and required the evaluation of five embedded integrals, as demonstrated in Appendices A and H. Importantly, we stress that in the present case, the LBD (equation 5) allowed us to analytically perform the angular part of the orbit average (equation 4). This helped reducing the computation time. We refer to Appendix C.1 for details.
@@
Figure 8: Relaxation rate, $\partial F/\partial t$, in ${(J_{\mathrm{r}},L)}$ as in Fig. 7, but here from the NR theory. The prediction matches the $N$-body measurements from Fig. 7, up to an overall prefactor that depends on $q$, and (very) weakly on $\alpha$.
@@
When comparing the $N$-body measurements (Fig. 7) with the NR prediction (Fig. 8), the NR theory successfully recovers the in-plane relaxation of stable rotating clusters: both figures exhibit strikingly similar structures in action space. We obtain results for the non-rotating case that are in agreement with T22. Yet, even though amplitudes in both approaches are comparable, there is still a (slight) overall prefactor mismatch. The inclusion of rotation has little impact on this mismatch. A more quantitative comparison would require performing (many) more $N$-body runs, as well as a more precise measurement of the relaxation rate. This would be no light undertaking, and will be the subject of future works.
@@ 5.1 $N$-body measurements
Let us now consider the rotating clusters presented in the middle and bottom rows of Fig. 9. In radially anisotropic (${q\!=\!1}$) and isotropic (${q\!=\!0}$) clusters, the systems lose stars in the prograde region (${\cos I\!>\!0}$) and gain stars in the retrograde region (${\cos I\!<\!0}$). As one increases $\alpha$ (i.e. as one increases the net rotation), this trend strengthens. This is in agreement with the orbital reshuffling observed in Fig. 8, where the clusters always tend to isotropise their in-plane distribution. Finally, we note that the highest diffusion rates in inclination are observed near ${\cos I\!=\!0}$. This corresponds precisely to the location of the discontinuity of the LBD.